Posts - cdr
Consolidated Data Return which is the collation of FELS, FESR, EF, ESER and FLU

#NICIS Managers meeting at #CNI. Main topics discussed were #CDR and the standardisation of curriculum definition in #QLS across #Allcolleges


#Change CR0106 issued to change board #BSUManagementGroup. This is for the removal of selected fields from the #CDR output file.


#Change CR0105 issued to change board #BSUManagementGroup. This is for the inclusion of postcode related data in the #CDR output file.


#Change CR0104 issued to change board #BSUManagementGroup. This is for the inclusion of non-QLS variables in the #CDR output file.


#Change CR0103 issued to change board #BSUManagementGroup. This is for the inclusion of equality fields in the #CDR output file.


Helped set up #CDR to be run locally @ #SRC


Working with #nwrc on running #cdr locally @ #allcolleges


Starting work on Validation set 2 for #CDR with view to have completed and signed off by end of week. #Allcolleges


Productive meeting at #DEL about #CDR and the plan to completion.


Project plan to completion of #CDR agreed with #BSU Management Group. Includes plan to allow local running of #CDR at #Allcolleges


Query from #BMC regarding #CDR validations


#CDR Validation query from #SWC


#CDR review meeting with #DEL in relation to weekend run.


Validation breakdown sent to #allcolleges from #cdr run carried out on Friday evening.


Successful run of #CDR on the 24th Jan. #allcolleges


meeting with #DEL regarding #CDR validation reconciliation.


Request from #NWRC to provide #CDR table script.


Produce paper on future support of #CDR and #FLU for consideration by #DEL and #Allcolleges


Investgation into failure notifications within the repository server. #CDR #Allcolleges


Analysing data from Monday evenings #CDR run.#All colleges


#DEL want a list of DEL employees that are currently set up to receive #CDR outputs.


#CDR set up for #test run this evening. #allcolleges


#change CR0087 sent to BSU management group for aproval . #CDR


Request from #BMC to provide #CDR run


Update continuing to #cdr #Validations


Update continuing to #cdr #Validations


Request from #SRC to reset password to enable access to #CDR files.


#CDR run completed on the 15th November.


Request from #BMC to send #CDR files for academic years, 10/11,11/12,12/13,13/14.


Updates to be made to #CDR #Validations based on weekend run. #allcolleges,#DEL


Creation of Look up tables for #CDR #Validations. #Colleges


Papers created for Thursdays #CDR meeting, showing results from SPSS and Sql #Validations. #allcolleges


Meeting arranged for Thursday the 24th to discuss #CDR validations. #allcolleges


Weekend run of #CDR for academic year 13/14. Run went ok just waiting feed back from #DEL regarding data. #allcolleges


Testing of the #CDR with #Validations that took place at the weekend.


1st of #CDR #Validations recieved. Going to merge and test these with already existing ones. #allcolleges


#CDR validations distributed to #Allcolleges for development. This will complete phase 2 of the project thus allowing all statistical returns to be made to #DEL from #CDR from 2013-14


#CDR #Validation guide to be created for #allcolleges. #DEL


#Change CR0082 accepted. #CDR changes to be carried out ASAP. #Allcolleges


Detailed meeting with #DEL and #CDR project group to agree the final list of validations and timescales for phase 2 of the project. #Allcolleges will be involved in assisting with validation development from 9 September 2013


Meeting to discuss project plan and way forward with validations to enable #CDR to be used for 2013/14 returns in November.


#CDR #Validation discussion meeting with #NWRC. Taking a look at Validations suggested by #DEL in prepartion for #DEL meeting.


Asked by #DEL to investigate #CDR differnces withing #SERC and #SWC


Results from #CDR test from the 25th of July have been recieved from #DEL. Anomolies to be looked into. #allcolleges


#BSUGroup meeting at #CNI. Agreed proposal for #Allcollege involvement in #CDR phase 2 developments and #QLP process harmonisation


Testing off #CDR in #NWRC to see if the #FESRrebuild is causing problems


List of differnces from last #CDR to be investigated. #DEL. #AllColleges


#CDR meeting to analyse results from previous run. #AllColleges


#CDR review from run the took place in May. Agreement to run again to see if issues/discrepancies have been reduced. #AllColleges.


Discussion with #BMC #Curriculum with regards to discrepancies being produced from the #CDR


Investigation into discrepancies with #CDR and #NRC data.


Meeting with #DEL regarding the handover of the #CDR project to myself. Ongoing issues and to do's where dicussed and also where we were sitting with regards to completion of phase 1. #AllColleges


Change request CR0078 for amendments to the #CDR output has been implemented and tested. #DEL #BMC #NRC #NWRC #SERC #SRC #SWC


Issuing a change request for some improvements and amendments to the #CDR extract....testing on-going, all good so far.


Meeting with #DEL to review #CDR and use of end dates going forward. Continuing testing and consolidation against individual returns. Positive results so far.


Investigating end dates in #EssentialSkills for #DEL as part of the #CDR project.


Testing #CDR variables with #NWRC.


Distribute #BSUGroup minutes of the 19 February 2013 meeting to all members. Key decisions included #SSRS training for #BSU staff and #NICISswg, college testing of #CDR and launch of 'this #NICIS hub'.


Gentle reminder to #Mhance for copy of filters applied to the Essential Skills return to further the testing of the #CDR project.


#CDR review meeting today with #DEL and #NICISswg Chairperson. Agreed to collate all current return specifications and review for consistency over next 2 weeks. Further review meeting at that stage.


Initial responses from #DEL on #CDR testing indicate positive looking results #BMC #NRC #NWRC #SERC #SRC #SWC


#FLU and #CDR ran successfully from the datacentre for all colleges over the weekend #BMC #NRC #NWRC #SERC #SRC #SWC #DEL


Meeting with #DEL and #NICISswg representative to discuss next steps with #CDR before review meeting on 19 February. Need to include #FTE details and other flags to enable more effective use of return.


#CDR extract changes - colleges to receive output 8th Feb lift #BMC #NRC #NWRC #SERC #SRC #SWC #DEL


#ChangeRequest CR0075 to allow #SRC #BMC #SERC #SRC #SWC and #NWRC access to #CDR output file approved


#CDR extract changes - colleges to receive output 8th Feb lift #BMC #NRC #NWRC #SERC #SRC #SWC #DEL


Agree with #DEL and #BSUgroup that #CDR cannot use current method of withdawn identifiaction used in #FLU at present


Cancel #ChangeRequest CR0074 and replace with CR0075 to allow college access to #CDR output only


Arrange #CDR phase 1 project review meeting with #DEL - to take place on 19 February 2013


#NICISswg agenda items includes #CDR #MISOBC and ideas for sharing ideas and functionality across colleges


Initial #ChangeRequest CR0074 to allow colleges access to #CDR data files plus add add flags for retention (based on same rule as #FLU


Attend #BSUGroup meeting where #FLU and #CDR support going forward was discussed along with #MISOBC and current #NICIS contract with #M-Hance


#BSUgroup meeting discussing #CDR #MISOBC and #Reporting Strategy


Supplier quarterly support meeting with #M-Hance. Discuss #QLP upgrades, #CDR support, #FLU changes for 2013, Support calls made by #SRC #BMC #SERC #SRC #SWC and #NWRC


Attend #NICISswg where PWC were inattendance discussing the #MISOBC requirements. Aldo #CDR overview provided


Meet #DEL to review progress with #CDR and plot next steps


#CDR meeting to review results from recent test


#CDR testing with #DEL going well


#CDR testing with #DEL going well


#CDR review meeting with #DEL